Please can you help me in this question because am having alot of problem...
there it says complete the table of values for y=3-x

Answer by praseenakos@yahoo.com(507)

Question:

complete the table of values for y=3-x


Answer:

As per the given question, you have to find the values of y corresponding to some values of x.

here you have given that, y = 3-x

Now give some values for x. (If you are given some values for x with the question, then you can use those values....otherwise take some arbitrary values...)

When x = 0, y = 3 -0 = 3
When x = 1, y = 3 - 1 = 2
When x = 2, y = 3 - 2 = 1
when x = 3, y = 3 - 3 = 0


Like wise you can find out the values of y corresponding to the values of x.


Hope you found the explanation useful.

Regards.

Praseena.